# Label-Efficient Cross-Modality Generalization for Liver Segmentation in Multi-Phase MRI

### v1

@@ 1 Introduction
Medical image segmentation has existed and been an important task for a long time, primarily due to its inherent interpretable nature, which is helpful to doctors in early diagnosis and treatment [13, 1, 15]. To date, there have been numerous medical image segmentation studies that were conducted on various organs of the human body, and one of them is the liver [13, 22, 9, 6]. Especially, under the label-scarce scenario, many noticeable works[17, 18, 16] have tackled the partially annotated medical datasets and shown promising results on various settings.
The liver is one of the most important internal organs responsible for critical physiological functions. Additionally, the liver has complex anatomical structures [10]. This complexity is further exacerbated by coupling with various pathological changes and high heterogeneity in medical images. One of the common liver diseases with such pathological changes today is liver fibrosis, a progressive condition resulting from chronic liver injury. Therefore, to diagnose effectively, it is crucial to segment the liver and accurately assess the fibrosis stage.
In response to these clinical needs, the CARE (Comprehensive Analysis & computing of REal-world medical images) 2025 challenge [12, 5, 21], in conjunction with MICCAI 2025, aims to advance real-world medical image analysis. In the CARE 2025 challenge, there are a total of five tracks that focus on distinct organs and their specific problems. In these tracks, CARE-Liver aims to advance the progress of Liver Fibrosis quantification and analysis methods. This track includes a crucial task, LiSeg - the automatic liver segmentation for liver MRI scans from multi-sequence and multi-center with limited ground truth.
@@
In this paper, we propose a label-efficient liver segmentation framework tailored for multi-phase, multi-vendor MRI under limited annotation and cross-modality domain shift. To address the scarcity of annotated hepatobiliary-phase (GED4) data and the distributional gap across imaging centers, we leverage a suite of semi-supervised learning techniques, including Cross Pseudo Supervision (CPS)[3], Bidirectional Copy-Paste (BCP)[2], and MiDSS[14], to exploit unlabeled GED4 and non-contrast sequences effectively. As our backbone, we adopt STU-Net[7], a scalable and transferable segmentation model pretrained on large-scale multi-organ datasets, and further fine-tune it on the ATLAS liver segmentation dataset to improve adaptation to the target domain. This fine-tuning step provides a strong initialization prior to applying semi-supervised strategies. We validate our approach on the CARE-Liver challenge LiSeg task and conduct extensive ablation studies to analyze the contribution of each component in our pipeline, demonstrating strong generalization across MRI modalities and scanner vendors with minimal supervision.
@@ 2.2 Scalable and Transferable Medical Image Segmentation Models
Although nnU-Net [8] has demonstrated strong performance across various tasks by automatically adapting to dataset-specific properties (e.g., input patch size, spacing), its reliance on dynamic hyperparameter selection limits the transferability of pretrained models. Core architectural parameters such as the number of resolution stages, kernel sizes, and downsampling/upsampling ratios are derived from dataset heuristics, making nnU-Net models less generalizable across tasks without retraining.
To address this, STU-Net [7] introduces a scalable and transferable framework by categorizing hyperparameters into two types: (i) weight-related (e.g., resolution stages), and (ii) weight-unrelated (e.g., input patch size). The former are fixed (e.g., six resolution stages and $(3\times 3\times 3)$ convolution kernels), ensuring consistent model weights across tasks, while the latter follow nnU-Net’s robust defaults.
STU-Net further incorporates several architectural improvements: (1) Residual blocks replace basic Conv-IN-LeakyReLU blocks in the encoder and decoder, using two $3\times 3\times 3$ convolutions with a skip connection to stabilize deep training and preserve gradient flow. (2) Downsampling blocks employ a dual-branch design: a main branch with two convolutions (stride 1 then 2) and a skip branch with a $1\times 1\times 1$ convolution (stride 2), and their outputs are summed to retain spatial features. (3) Upsampling blocks use nearest-neighbor interpolation followed by a $1\times 1\times 1$ convolution, avoiding learnable parameters and improving transferability across resolutions. (4) Compound scaling, inspired by EfficientNet [20], jointly increases network depth and width at each resolution stage, enhancing both receptive field and representational capacity.
@@ 3.1 Dataset
LiQA Dataset The dataset comprises a multi-center cohort of 610 patients diagnosed with liver fibrosis, including 170 new cases beyond the CARE2024 challenge [12, 5, 21]. Patients were scanned across four clinical centers using three MRI vendors Philips Ingenia 3.0T, Siemens Skyra 3.0T, and Siemens Aera 1.5T, resulting in realistic multi-vendor variability. Each patient underwent multi-phase MRI, including T2-weighted (T2WI), diffusion-weighted (DWI), T1-weighted (T1WI), and Gd-EOB-DTPA–enhanced dynamic imaging at four phases: arterial (GED1, 25s post-injection), portal venous (GED2, 1min), delayed (GED3, 4min), and hepatobiliary (GED4, 20min). The training set includes 360 cases (vendor A: 130, B1: 170, B2: 60), but only 30 GED4 scans (8.3%) have segmentation labels, posing a strong annotation scarcity challenge. Consequently, subtask 1 (T1WI, T2WI, DWI) operates fully unsupervised, while subtask 2 (GED4) requires label-efficient or semi-supervised strategies. The validation set contains 60 fully annotated cases (20 per vendor A, B1, B2). The test set includes 190 cases from all vendors (A: 40, B1: 40, B2: 40, C: 70), with vendor C held out to assess generalization. Key challenges include missing sequences (except GED4), inter-vendor protocol differences, and the lack of spatial pre-alignment, reflecting the heterogeneity of clinical liver MRI data.
ATLAS Dataset The ATLAS dataset from the MICCAI 2023 ATLAS challenge [19], includes 60 public and 30 private patient cases. Each case contains multiple NIFTI-format MRI scans and corresponding labels. Images were acquired using various MRI machines and sequences, with one of three post-contrast phases (arterial, portal, or delayed) selected. Liver and tumour masks were merged into a single foreground for the accurate liver segmentation task.
@@ 3.3 Results
Sub-Task 1: Contrast-Enhanced Liver Segmentation (GED4) The table 2 shows the results on contrast-enhanced hepatobiliary phase (GED4) using the private validation set. Three methods compared include the state-of-the-art semi-supervised model CPS [3], BCP [2], and MiDSS [14]. Specifically, CPS achieves superior performance with a DSC of 0.9685 and an HD of 25.56. Both BCP and MiDSS, which adopt teacher–student architectures, achieved competitive results but still underperform compared to CPS.
@@
Regarding the different features, physically, T2WI exhibits completely reversed signal characteristics compared to GED4 and T1WI. Specifically, while the liver appears hyperintense (bright) on GED4 due to hepatocytes’ uptake of Gd-EOB-DTPA intermediate-to-high signal on T1WI, the liver appears hypointense (dark) on T2WI [4]. This contrast pattern reversal creates a fundamental challenge for CNN feature transfer [11]. CNN, which learns from 30 annotated GED4 images, tends to detect bright-to-dark transitions such as liver-to-background. In contrast, when it is applied to T2WI with dark liver regions, its features fail catastrophically with significant performance drops for the cross-modality medical image segmentation task.

### v3

@@ 1 Introduction
Medical image segmentation has existed and been an important task for a long time, primarily due to its inherent interpretable nature, which is helpful to doctors in early diagnosis and treatment [ma2024segment, azad2024medical, nguyen2023towards]. To date, there have been numerous medical image segmentation studies that were conducted on various organs of the human body, and one of them is the liver [ma2024segment, zhu2024medical, isensee2024nnu, heller2023kits21]. Especially, under the label-scarce scenario, many noticeable works[nguyen2025semi, pham2025fetal, nguyensemi] have tackled the partially annotated medical datasets and shown promising results on various settings.
The liver is one of the most important internal organs responsible for critical physiological functions. Additionally, the liver has complex anatomical structures [juza2014clinical]. This complexity is further exacerbated by coupling with various pathological changes and high heterogeneity in medical images. One of the common liver diseases with such pathological changes today is liver fibrosis, a progressive condition resulting from chronic liver injury. Therefore, to diagnose effectively, it is crucial to segment the liver and accurately assess the fibrosis stage.
In response to these clinical needs, the CARE (Comprehensive Analysis & computing of REal-world medical images) 2025 challenge [liu2025merit, gao2023reliable, wu2022meru], in conjunction with MICCAI 2025, aims to advance real-world medical image analysis. In the CARE 2025 challenge, there are a total of five tracks that focus on distinct organs and their specific problems. In these tracks, CARE-Liver aims to advance the progress of Liver Fibrosis quantification and analysis methods. This track includes a crucial task, LiSeg - the automatic liver segmentation for liver MRI scans from multi-sequence and multi-center with limited ground truth.
@@
In this paper, we propose a label-efficient liver segmentation framework tailored for multi-phase, multi-vendor MRI under limited annotation and cross-modality domain shift. To address the scarcity of annotated hepatobiliary-phase (GED4) data and the distributional gap across imaging centers, we leverage a suite of semi-supervised learning techniques, including Cross Pseudo Supervision (CPS)[chen2021semi], Bidirectional Copy-Paste (BCP)[bai2023bidirectional], and MiDSS[ma2024constructing], to exploit unlabeled GED4 and non-contrast sequences effectively. As our backbone, we adopt STU-Net[huang2023stu], a scalable and transferable segmentation model pretrained on large-scale multi-organ datasets, and further fine-tune it on the ATLAS liver segmentation dataset to improve adaptation to the target domain. This fine-tuning step provides a strong initialization prior to applying semi-supervised strategies. We validate our approach on the CARE-Liver challenge LiSeg task and conduct extensive ablation studies to analyze the contribution of each component in our pipeline, demonstrating strong generalization across MRI modalities and scanner vendors with minimal supervision.
@@ 2.2 Scalable and Transferable Medical Image Segmentation Models
Although nnU-Net [isensee2021nnu] has demonstrated strong performance across various tasks by automatically adapting to dataset-specific properties (e.g., input patch size, spacing), its reliance on dynamic hyperparameter selection limits the transferability of pretrained models. Core architectural parameters such as the number of resolution stages, kernel sizes, and downsampling/upsampling ratios are derived from dataset heuristics, making nnU-Net models less generalizable across tasks without retraining.
To address this, STU-Net [huang2023stu] introduces a scalable and transferable framework by categorizing hyperparameters into two types: (i) weight-related (e.g., resolution stages), and (ii) weight-unrelated (e.g., input patch size). The former are fixed (e.g., six resolution stages and $(3\times 3\times 3)$ convolution kernels), ensuring consistent model weights across tasks, while the latter follow nnU-Net’s robust defaults.
STU-Net further incorporates several architectural improvements: (1) Residual blocks replace basic Conv-IN-LeakyReLU blocks in the encoder and decoder, using two $3\times 3\times 3$ convolutions with a skip connection to stabilize deep training and preserve gradient flow. (2) Downsampling blocks employ a dual-branch design: a main branch with two convolutions (stride 1 then 2) and a skip branch with a $1\times 1\times 1$ convolution (stride 2), and their outputs are summed to retain spatial features. (3) Upsampling blocks use nearest-neighbor interpolation followed by a $1\times 1\times 1$ convolution, avoiding learnable parameters and improving transferability across resolutions. (4) Compound scaling, inspired by EfficientNet [tan2019rethinking], jointly increases network depth and width at each resolution stage, enhancing both receptive field and representational capacity.
@@ 3.1 Dataset
LiQA Dataset The dataset comprises a multi-center cohort of 610 patients diagnosed with liver fibrosis, including 170 new cases beyond the CARE2024 challenge [liu2025merit, gao2023reliable, wu2022meru]. Patients were scanned across four clinical centers using three MRI vendors Philips Ingenia 3.0T, Siemens Skyra 3.0T, and Siemens Aera 1.5T, resulting in realistic multi-vendor variability. Each patient underwent multi-phase MRI, including T2-weighted (T2WI), diffusion-weighted (DWI), T1-weighted (T1WI), and Gd-EOB-DTPA–enhanced dynamic imaging at four phases: arterial (GED1, 25s post-injection), portal venous (GED2, 1min), delayed (GED3, 4min), and hepatobiliary (GED4, 20min). The training set includes 360 cases (vendor A: 130, B1: 170, B2: 60), but only 30 GED4 scans (8.3%) have segmentation labels, posing a strong annotation scarcity challenge. Consequently, subtask 1 (T1WI, T2WI, DWI) operates fully unsupervised, while subtask 2 (GED4) requires label-efficient or semi-supervised strategies. The validation set contains 60 fully annotated cases (20 per vendor A, B1, B2). The test set includes 190 cases from all vendors (A: 40, B1: 40, B2: 40, C: 70), with vendor C held out to assess generalization. Key challenges include missing sequences (except GED4), inter-vendor protocol differences, and the lack of spatial pre-alignment, reflecting the heterogeneity of clinical liver MRI data.
ATLAS Dataset The ATLAS dataset from the MICCAI 2023 ATLAS challenge [quinton2023tumour], includes 60 public and 30 private patient cases. Each case contains multiple NIFTI-format MRI scans and corresponding labels. Images were acquired using various MRI machines and sequences, with one of three post-contrast phases (arterial, portal, or delayed) selected. Liver and tumour masks were merged into a single foreground for the accurate liver segmentation task.
@@ 3.3 Results
Sub-Task 1: Contrast-Enhanced Liver Segmentation (GED4) The table 2 shows the results on contrast-enhanced hepatobiliary phase (GED4) using the private validation set. Three methods compared include the state-of-the-art semi-supervised model CPS [chen2021semi], BCP [bai2023bidirectional], and MiDSS [ma2024constructing]. Specifically, CPS achieves superior performance with a DSC of 0.9685 and an HD of 25.56. Both BCP and MiDSS, which adopt teacher–student architectures, achieved competitive results but still underperform compared to CPS.
@@
Regarding the different features, physically, T2WI exhibits completely reversed signal characteristics compared to GED4 and T1WI. Specifically, while the liver appears hyperintense (bright) on GED4 due to hepatocytes’ uptake of Gd-EOB-DTPA intermediate-to-high signal on T1WI, the liver appears hypointense (dark) on T2WI [curvo2010hypointense]. This contrast pattern reversal creates a fundamental challenge for CNN feature transfer [kang2024target]. CNN, which learns from 30 annotated GED4 images, tends to detect bright-to-dark transitions such as liver-to-background. In contrast, when it is applied to T2WI with dark liver regions, its features fail catastrophically with significant performance drops for the cross-modality medical image segmentation task.